Lars:

Solve the system using elimination. (Show your work) 3x + y = 10 y = 6x + 1

Lars:

Why'd it post this question again bro 😭😭

jhonyy9:

so like a first step re arrange the second equation to be the same like the first one 3x +y = 10 -6x +y = 1 - in the next step multiplie the first equation or the second equation by -1 and add (assume) the same terms 3x +y = 10 I × (-1) -3x -y = -10 -6x +y = 1 -3x + (-6x) + (-y) +y = -10 +1 -9x +0 = -9 -9x = -9 divide both sides by -9 x = 1 hope from here you can calcule easy the y value